### v1

# On the excitability of two-level atoms by spectrally encoded single-photon wave packets

Abstract
Single-photon wave packets are utilized as quantum information carriers within quantum information science and quantum communication applications. Here, we analyze the time-dependent interaction of a two-level atom with a spectrally encoded single-photon wave packet and obtain the excitation probability of the atom based on the Heisenberg-Langevin equations. Our results show that the spectral encoding process causes the single-photon wave packet to spread in the time domain, and the intensity of the wave packet decreases. Therefore, an encoded single-photon wave packet excites a two-level atom much weaker than an uncoded one. These results are crucial for understanding how encoded quantum light interacts with quantum nodes in realistic quantum networks. In particular, they reveal limitations and trade-offs in using spectrally encoded photons for secure, multiplexed quantum communication. The findings offer insights into optimizing encoding strategies for efficient atom-photon interactions and contribute to the development of scalable and secure quantum networking protocols.
keywords: Atom-photon interaction, Heisenberg-Langevin equations, Spectral encoding, Photon wave packet.
I Introduction
In recent years, much attention has been paid to the interaction between atom and light, which has many applications in quantum information, quantum communication, and fundamental experiments in quantum mechanics Lewenstein ; Rui ; Andersson ; Han ; Diaz ; Shomroni ; Tirandaz ; Moeini . In this regard, a single-photon wave packet is one of the most important candidates for examining atom-photon interaction in the experiments of quantum information processing Furusawa ; Brecht . Although quantum information can be encoded in different physical systems, such as trapped ions and superconducting devices, photonic states are much more practical because photons are indeed more robust against decoherence than many other quantum systems Flamini . Their ability to travel through optical fibers and free space with minimal loss makes them practical for carrying information between various nodes of a quantum network Gheri . Therefore, single-photon wave packets are good candidates for researchers in many fields, especially in quantum communication Hu , quantum computation Sheng2 , quantum key distribution Sych1 ; Sych2 , and entanglement Sheng1 .
Recently, many studies have examined the interaction between photon wave packets and atoms Slodicka ; Aljunid , molecules Wrigge , surface plasmons Chang and quantum dots Vamivakas . Furthermore, the interaction between a two-level atom and a single-photon wave packet has been extensively analyzed Lindlein ; Rephaeli ; Chen ; Rag ; Wang1 . However, the interaction of a two-level atom and an encoded single-photon wave packet has not been studied much.
In this paper, we analyze the time-dependent interaction between a two-level atom with an encoded single-photon wave packet to find the excitation probability of the atom which is in the ground state, based on the Heisenberg-Langevin equations. Then, we compare our results with the typical dynamic of interaction between a two-level atom and a single-photon wave packet reported in the literature Domokos ; Wang2 ; Tannoudji . We use the binary spectral encoding approach to generate the encoded single-photon wave packet provided in code division multiple-access (CDMA) communication systems Salehi .
The paper is organized as follows. In section II, we review a general formalism to analyze the dynamics of a two-level atom that interacts with a continuous mode state of quantum light. In section III, first, we introduce the spectral encoding approach. Then, we investigate the dynamics of interaction between a two-level atom with an encoded single-photon wave packet to find the excitation probability of the atom. Finally, the results are discussed in section IV.
II Model
A general formalism to study the dynamics of a two-level atom that interacts with a continuous mode state of quantum light is presented in Domokos ; Wang2 .
Let us consider the interaction of a two-level atom with the quantized photon wave packet as depicted in FIG. 1. The continuous mode state of a quantized field operator in the interaction picture is Loudon
$E(t)=i\int\operatorname{d}\omega(\frac{\hbar\omega}{4\pi\epsilon_{0}cA})^{1/2}\big{[}a(\omega)e^{-i\omega t}-a^{\dagger}(\omega)e^{i\omega t}\big{]}$ (1)
where $a^{\dagger}(\omega)$ and $a(\omega)$ are the creation and annihilation operators of the quantized field with the frequency $\omega$, respectively. The commutation relation of these operators is $\big{[}a(\omega),a^{\dagger}(\omega)\big{]}=\delta(\omega-\omega^{\prime})$, where $\delta$ denotes the Dirac delta function. Moreover, $\epsilon_{0}$, $c$ and $A$ are the vacuum permittivity, the speed of light and the cross section, respectively.
Figure 1: A schematic of the interaction between a two-level atom and a photon wave packet.
In the interaction picture and using rotating wave approximation (RWA), one can show that the interaction Hamiltonian of the system can be written as
$H_{\operatorname{I}}=-i\hbar\int\operatorname{d}\omega\lambda_{\omega}\big{[}\sigma_{+}a(\omega)e^{-i(\omega-\omega_{A})t}-\operatorname{H.C}\big{]}$ (2)
where $\omega_{A}$ is the atomic transition frequency and $\sigma_{z}=|e\rangle\langle e|-|g\rangle\langle g|$, $\sigma_{+}=|e\rangle\langle g|$, $\sigma_{-}=|g\rangle\langle e|$ are the atomic operators and H.C denotes Hermitian conjugates. Furthermore, the coupling strength is $\lambda_{\omega}=d\big{(}\omega/4\pi\epsilon_{0}\hbar cA\big{)}^{1/2}$ where $d$ is the value of the dipole moment of the atom.
One can show that the evolution of the atomic operators can be described by a set of Heisenberg-Langevin equations as Domokos ; Wang2 ; Tannoudji
$\frac{\operatorname{d}\sigma_{-}}{\operatorname{d}t}=-\gamma\sigma_{-}/2+\sigma_{z}\int\operatorname{d}\omega\lambda_{\omega}a(\omega)e^{-i(\omega-\omega_{A})t}+\zeta_{-}$ (3)
$\frac{\operatorname{d}\sigma_{z}}{\operatorname{d}t}=-\gamma(\sigma_{z}+1)-2\int\operatorname{d}\omega\lambda_{\omega}\big{[}\sigma_{+}a(\omega)e^{-i(\omega-\omega_{A})t}+\operatorname{H.C}\big{]}+\zeta_{z}$ (4)
where $\gamma$ is the decay rate of the atomic dipole and can be defined as $\gamma=2\pi\int\operatorname{d}\omega|\lambda_{\omega}|^{2}\delta(\omega-\omega_{A})$. Since we also considered the interaction between the atom and the environment, the terms of noise operators ($\zeta_{z}$ and $\zeta_{-}$) have appeared in the Heisenberg-Langevin equations, which are defined as
$\zeta_{-}=\sigma_{z}\int\operatorname{d}\omega^{\prime}\lambda_{\omega^{\prime}}b(\omega^{\prime})e^{-i(\omega^{\prime}-\omega_{A})t}$ (5)
$\zeta_{z}=$ $-2\int\operatorname{d}\omega^{\prime}\lambda_{\omega^{\prime}}\big{[}\sigma_{+}b(\omega^{\prime})e^{-i(\omega^{\prime}-\omega_{A})t}+\operatorname{H.C}\big{]}$ (6)
where $b^{\dagger}(\omega^{\prime})$ and $b(\omega^{\prime})$ are the creation and the annihilation operators of the environmental modes with the frequency $\omega^{\prime}$, respectively.
From now on, for simplicity, we use the narrowband approximation means that the spectral width of photon-wavepacket $\xi(\omega)$ is small compared to its central frequency $\omega_{0}$ Loudon . So, the coupling strength can be written as $\lambda_{\omega}=d\big{(}\omega_{0}/4\pi\epsilon_{0}\hbar cA\big{)}^{1/2}$.
II.1 Excitation of a two-level atom by a rectangular single-photon wave packet
We now calculate the excitation probability $P_{e}(t)$ of a two-level atom in interaction with a single-photon wave packet which can be defined as Scully
$P_{e}(t)=\frac{1}{2}\big{[}\langle\Psi_{0}|\sigma_{z}(t)|\Psi_{0}\rangle+1\big{]}$ (7)
We suppose the initial state of the total system is $|\Psi_{0}\rangle=|g\rangle|1_{\xi}\rangle|0_{e}\rangle$ where is a product state of the atomic ground state, the single-photon wave packet state, and the vacuum state of the environment, respectively.
Figure 2: The excitation probability $P_{e}(t)$ of a two-level atom that interacts with a rectangular shape of a single-photon wave packet as a function of $\gamma t$.
A single-photon wavepacket state can be defined as $|1_{\xi}\rangle=a^{\dagger}_{\xi}|0\rangle$ where
$a^{\dagger}_{\xi}=\int\operatorname{d}\omega\xi(\omega)a^{\dagger}(\omega)$ (8)
where $\xi(\omega)$ and $a^{\dagger}_{\xi}$ represent the spectral amplitude of the photon wave packet and the photon wave packet creation operator, respectively Loudon . Moreover, we assume the environment to be initially in the vacuum state. Based on the Born approximation, the changes in the state of the environment are negligible Schlosshauer . So, the state of the environment remains vacuum at all times. Considering the initial vacuum state of the environment, the average values of the noise operators in Eq. (3) and Eq.(4) will be zero.
To obtain the excitation probability $P_{e}(t)$ of a two-level atom, we need to calculate the expectation value of the atomic operator $\sigma_{z}$ as $\langle\sigma_{z}(t)\rangle$ in Eq. (7). So, one should solve the set of coupled differential equations for the atomic operators ($\sigma_{z}$ and $\sigma_{\pm}$) in Eq. (3) and Eq. (4). Moreover, we should define the explicit shape of the single-photon wave packet.
In this regard, we take a single-photon wave packet with a rectangular shape as an example Salehi :
$\xi(\omega)=\begin{cases}\sqrt{P_{0}}/W&\mbox{for}-\frac{W}{2}\leq\omega\leq\frac{W}{2}\\ 0&\mbox{elsewhere}\end{cases}$ (9)
where $P_{0}$ and $W$ are the peak power and the total bandwidth of the pulse, respectively. Therefore, one can show that the temporal shape pulse is a sinc function:
$\xi(t)=\sqrt{P_{0}}\mbox{sinc}(\frac{W}{2}t)$ (10)
where sinc x=sin $\mbox{sinc x}=\frac{\mbox{sin x}}{\mbox{x}}$.
Assuming $W=1.5\gamma$, we can obtain the excitation probability $P_{e}(t)$ of a two-level atom that interacts with the rectangular shape of a single-photon wave packet. In FIG. 2, we have plotted the excitation probability $P_{e}(t)$ of the atom as a function of $\gamma t$. This figure shows the damping behavior of the atom in the presence of the environment.
III Encoded single-photon wave packets
In this section, we study the interaction between a two-level atom and an encoded single-photon wave packet. First, we review the spectral phase encoding approach.
III.1 Spectral encoding approach
The spectral phase encoding approach to encode the spectrum of a photon wave packet is used in CDMA communication systems. See Salehi for a general review.
A spectral phase-shifting operator $U$ with a phase value $\theta$ can be written as
$U=e^{-i\sum_{\omega}\theta(\omega)a^{\dagger}(\omega)a(\omega)}=\prod_{\omega}U(\omega)$ (11)
where $\theta(\omega)$ is the phase shift of the frequency $\omega$ on the spectral wave packet of the pulse $\xi(\omega)$ and $U(\omega)=e^{-i\theta(\omega)a^{\dagger}(\omega)a(\omega)}$. In the Heisenberg picture, one can show that:
$Ua^{\dagger}(\omega)U^{\dagger}=a^{\dagger}(\omega)e^{-i\theta(\omega)}$ (12)
Applying Eq. (12) to the photon wave packet creation operator in Eq. (8), we have $Ua^{\dagger}_{\xi}U^{\dagger}=a^{\dagger}_{\xi^{e}}$ where
$\xi^{e}(\omega)=\xi(\omega)e^{-i\theta(\omega)}$ (13)
in which superscript $e$ denotes the encoded photon wave packet. Moreover, the encoded wave packet in the temporal domain can be obtained by the Fourier transform as
$\xi^{e}(t)=\frac{1}{\sqrt{2\pi}}\int\operatorname{d}\omega\xi(\omega)e^{-i\big{(}\omega t+\theta(\omega)\big{)}}$ (14)
Therefore, we conclude that the spectral encoding process just affects the spectral amplitude $\xi(\omega)$ of the quantum wave packet Salehi .
III.2 Excitation of a two-level atom by an encoded single-photon wave packet
To encode the single-photon wave packet with a rectangular shape in Eq. (10), we divided the spectrum of the wave packet $\xi(\omega)$ into $N_{0}$ sequential and distinct frequency chips, each bandwidth is $\Omega=W/N_{0}$. Based on the spectral encoding approach, one can show that the time domain representation of the encoded rectangular wave packet can be written as Salehi
$\xi^{e}(t)=\frac{\sqrt{P_{0}}}{N_{0}}\mbox{sinc}\big{(}\frac{\Omega}{2}t\big{)}\sum_{n=-N}^{N}\mbox{exp}\big{[}-i(n\Omega t)+\phi_{n}\big{]}$ (15)
where the phase $\phi_{n}$ denotes the $n$th element of a code consisting of $N_{0}=2N+1$. Moreover, the term $\mbox{sinc}\big{(}\frac{\Omega}{2}t\big{)}$ is a real function that represents the temporal width of the encoded wave packet and the exponential term shows a periodic behavior with period $T=2\pi/\Omega$ which its temporal shape depends on the code elements $\phi_{n}$. Note that if we have $\phi_{n}=0$ for all $n$, Eq. (15) reduces to Eq. (10) Salehi .
We now consider $N_{0}=3,5,7,31,63$ for the code length of an encoded single-photon wave packet to calculate the excitation probability $P_{e}(t)$ of a two-level atom which interacts with an encoded pulse according to Eq. (7). In addition, we assume the value of the phase $\phi_{n}$ is 00 and $\pi$ which correspond to the amplitude of $+1$ and $-1$, respectively. This spectral encoding process is called random binary spectral encoding Salehi .
We have plotted the excitation probability $P_{e}(t)$ of a two-level atom for the various code lengths of the encoded single-photon wave packet, which is compared to the uncoded wave packet in FIG. 3. We conclude that the encoding process of the single-photon wave packet decreases the excitation probability of a two-level atom. To analyze this result, we have plotted the intensity of the uncoded and encoded wave packet in FIG. 4. As is clear in this figure, the encoding process causes the single-photon wave packet to spread in time domain. Consequently, the power of the wave packet diminishes, and the wave packet behaves as a low-intensity pulse. Therefore, an encoded wave packet excites a two-level atom much weaker than an uncoded one.
Figure 3: The excitation probability $P_{e}(t)$ of a two-level atom which interacts with an encoded rectangular shape of photon wave packet as a function of $\gamma t$ for the various code length $N_{0}$ (blue curves) compared to the uncoded case (red curves). a) $N_{0}=3$ b) $N_{0}=5$ c) $N_{0}=7$ d) $N_{0}$=31 and e) $N_{0}=63$.
Figure 4: Intensity of the uncoded (orange) and encoded (blue) single-photon wave packet. In this figure, $N_{0}=31$.
The question that arises here is: what is the origin of the oscillations in the excitation probability of the atom that interacts with an encoded photon wave packet?. To answer this question, we investigate the relationship between the excitation probability of the atom and the sequences of the random phase (0 and $\pi$) in the spectral encoding process according to Eq. (15). In FIG. 5, we have plotted the excitation probability of the atom considering three different sequences of the random phase for the code length $N_{0}=63$. Our results show that the oscillations in the excitation probability of the atom depend on the selected sequences of the random phase in the binary spectral encoding process.
Figure 5: The excitation probability $P_{e}(t)$ of a two-level atom by an encoded photon wave packet considering three different sequences of the random phase (0 and $\pi$) for the code length $N_{0}=63$ in the binary spectral encoding process.
We also analyze the effect of the code length of the encoded photon wave packet on the excitation probability of a two-level atom. We have plotted the relationship between the excitation probability $P_{e}(t)$ of the atom and the code length $N_{0}$ of the encoded single-photon wave packet in FIG. 6. This figure shows that the excitation probability decreases with increasing code length. Since the code length increases, the photon wave packet spreading rate increases significantly (see FIG. 7). Consequently, the ability of the encoded wave packet to excite the atom decreases.
Figure 6: The excitation probability $P_{e}(t)$ of a two-level atom as a function of the code length $N_{0}$ of the encoded single-photon wave packet.
Figure 7: Spreading the encoded photon wave packet in time domain for the various code lengths. a) $N_{0}=3$ b) $N_{0}=5$ c) $N_{0}=7$ d) $N_{0}$=31 and e) $N_{0}=63$.
IV Conclusion
We studied the time-dependent interaction of a two-level atom with a spectrally encoded single-photon wave packet based on the Heisenberg-Langevin equations. We used a spectral phase encoding method inspired by classical CDMA systems to encode quantum information in the frequency domain of the photon. Considering this formalism, we showed that for an encoded wave packet, the excitation probability of the atom is much lower than an unencoded one. On the other hand, the encoding process causes the wave packet to spread in the time domain and the ability to excite a two-level atom decreases. Moreover, we evaluated how encoding parameters—such as code length and phase randomness—affect atomic excitation probabilities.
These findings are particularly relevant for quantum networks, where atoms serve as quantum nodes and qubits are transmitted over long distances via single-photon wave packets. In such systems, an encoded single-photon wave packet carries quantum information to the receiver. If the receiver applies a decoding mask that is the conjugate of the original encoding mask, the photon’s spectral profile is correctly reconstructed, allowing it to interact effectively with the atomic node. However, if the decoding mask does not match the encoding mask, the wave packet fails to recover its original shape and remains a temporally spread, low-intensity pulse. As a result, it has minimal impact on the atomic excitation. This behavior provides a built-in filtering mechanism: atoms can act as selective photodetectors, responding only to correctly encoded photons while rejecting mismatched or unauthorized signals. Such properties are valuable for enhancing the security and robustness of quantum communication protocols.
Moreover, this research provides a fundamental insight into how spectral encoding, commonly used in CDMA systems, affects the quantum mechanical behavior of light-matter interaction. Understanding this effect is essential for developing robust quantum communication protocols that rely on encoding information in the frequency domain. Our findings indicate that while spectral encoding enhances information security and multiplexing capabilities, it also introduces trade-offs in terms of interaction efficiency with quantum nodes, which must be addressed in the system design. This study also opens the door to future research in optimizing spectral encoding schemes that balance encoding complexity and atom-photon interaction strength. In practical implementations of quantum repeaters or quantum routers, such trade-offs must be well-understood to ensure high fidelity and reliability of quantum information transfer. Furthermore, the methodology presented here can be extended to investigate other encoding formats, multi-level atoms, and more complex photonic structures, including entangled wave packets or time-bin encodings.
Ultimately, our work contributes to the theoretical foundation necessary for advancing quantum network architectures and can inform experimental setups aiming to test photon-mediated interactions at the quantum level with encoded information. The ability to predict and control atomic excitation probabilities in the presence of encoded quantum signals is a step toward realizing scalable and secure quantum communication infrastructures.
Acknowledgment
The author would like to thank Jawad Salehi, Mohammad Rezai, and Ali Soltanmanesh for their instructive comments and useful discussion.
Data Availability
The data that support the findings of this study are available on request from the corresponding author